OpenStudy (anonymous):

Write an equation for the line that is parallel to y = -9x + 12 and that passes through the point (-6,-10).

OpenStudy (anonymous):

the lines are parallel meaning the gradient wud be same y=mx+c y=-9x+c then put the value of x and y in the above equation (y=-9x+c) to find the value of c x=-6 y=-10 -10=-9(-6)+c c=-10-54 c=-64 so the equation should become y=-9x-64

OpenStudy (anonymous):

@mohammad.arqum Thank you so much this is so much easier than my teacher teaches me!

OpenStudy (anonymous):

No problem...